OpenStudy (anonymous):

a bomber flying upward at 53deg with vertical release a bomb at at altitude of 800 m. the bomb strike the ground 20sec after its release.find velocity of bomber at release time

OpenStudy (anonymous):

 problem illustration...

OpenStudy (anonymous):

From Projectile: \[y = v_o\sin \theta \times t - \frac{1}{2}gt^2\]where \(\theta = 90-53=37\) Just plug-in other given values, you can solve for \(v_o\) which is the velocity of the bomber at the time the bomb is release...